# RAP-SAM [Uncaptioned image]: Towards Real-Time All-Purpose Segment Anything

### v1

@@ Abstract
Advanced by transformer architecture, vision foundation models (VFMs) achieve remarkable progress in performance and generalization ability. Segment Anything Model (SAM) is one remarkable model that can achieve generalized segmentation. However, most VFMs cannot run in real-time, which makes it difficult to transfer them into several products. On the other hand, current real-time segmentation mainly has one purpose, such as semantic segmentation on the driving scene. We argue that diverse outputs are needed for real applications. Thus, this work explores a new real-time segmentation setting, named all-purpose segmentation in real-time, to transfer VFMs in real-time deployment. It contains three different tasks, including interactive segmentation, panoptic segmentation, and video segmentation. We aim to use one model to achieve the above tasks in real-time. We first benchmark several strong baselines. Then, we present Real-Time All Purpose SAM (RAP-SAM). It contains an efficient encoder and an efficient decoupled decoder to perform prompt-driven decoding. Moreover, we further explore different training strategies and tuning methods to boost co-training performance further. Our code and model are available at https://github.com/xushilin1/RAP-SAM/.
@@ 1 Introduction
Figure 1: We present real-time all-purpose segmentation to segment and recognize objects for image, video, and interactive inputs. In addition to benchmarking, we also propose a simple yet effective baseline, named RAP-SAM, which achieves the best accuracy and speed trade-off among three different tasks. Both real-time panoptic segmentation and video instance segmentation are shown at the bottom.
@@
Recent advancements in computer vision, propelled by transformer architectures [13, 2, 75], focus on developing a singular, large-scale model versatile enough for various tasks, including detection and segmentation, with a general-purpose design. Specifically, several studies [74, 80] adopting identical architectures have demonstrated superior performance compared to models designed exclusively for specific tasks. Furthermore, numerous studies [34, 30] investigate the foundation model design using extensive data. A notable example, the Segment Anything Model (SAM) [34], introduces an interactive model driven by visual prompts for versatile segmentation. Concurrently, for semantic-level segmentation, numerous studies [10, 93, 47, 46, 29] employing a universal design have surpassed previous models tailored for specific image and video segmentation tasks.
However, the majority of these studies face challenges in achieving real-time performance or compatibility with mobile devices, primarily due to their heavy encoders and cascaded decoders. This constraint significantly limits the applicability of these advanced architectures in practical vision applications, although they can achieve high performance. On the contrary, various studies explore real-time segmentation using distinctive designs. Yet, these efforts predominantly concentrate on singular application purposes, including autonomous driving scenes, the COCO dataset for panoptic segmentation (PS) [50], and real-time video instance/object segmentation (VIS) [81]. Currently, no studies investigate real-time, all-purpose segmentation, wherein a model is capable of performing universal segmentation tasks, encompassing image segmentation, video instance/video segmentation, and SAM-like interactive segmentation.
@@
Inspired by the aforementioned challenges, this study investigates the novel issue of real-time, all-purpose segmentation. We pose a critical inquiry: given the limited computational resources and model capacity, how can we develop an efficient, all-purpose segmentation model? This entails creating a single model capable of segmenting, tracking, and classifying each pixel in real-time, akin to performing interactive segmentation similar to SAM.
@@
Initially, we define this research problem by integrating existing benchmarks. Specifically, we use the COCO [50] and YouTube-VIS 2019 [81] datasets for joint co-training, employing identical hyper-parameters. Beyond the semantic queries used for semantic-level segmentation in prior research, our approach additionally incorporates visual prompt queries in SAM for interactive segmentation. However, there are still several challenges: 1), No previous works explore joint co-training for three different purposes. Thus, extensive experiments are needed to search for a better meta-architecture. 2), Interactive segmentation and semantic-level segmentation need better strategies to balance each other. 3), Under the real-time setting, several previous architectures [80, 42] for multi-task segmentation may not be practical.
@@
To address these challenges, we benchmark several recent works [85, 10, 25] in a real-time setting. Then, we initially investigate a range of meta-architectures suitable for all-purpose segmentation. Unlike previous works, we avoid cascaded architecture and only use the pyramid feature once when performing dynamic convolution. We explore four different architectures in the case of various decoder designs, where we adopt a single decoder with multiple queries. Via extensive experiments, we adopt pooling-based dynamic convolution to replace per-pixel cross-attention to achieve better accuracy and speed trade-off. To balance interactive and semantic-level segmentation, we present an effective dual adapter design that better adapts shared knowledge of the same decoder. Finally, we term our best model as Real-time All-Purpose SAM (RAP-SAM). To the best of our knowledge, it is the first real-time all-purpose segmentation model. Moreover, due to the semantic level and interactive level segmentation ability, our model can also create more new applications, such as interactive video segmentation. Extensive experiments show the effectiveness of our design. The main contributions of this work are:
@@
  • We introduce all-purpose segmentation, a multi-task segmentation that aims to segment objects for image, video, and interactive inputs in real-time.
@@
  • We present a simple yet fast baseline named RAP-SAM. It contains a lightweight feature extractor, a unified decoder, and two asymmetric adapters.
@@
  • Extensive experiments demonstrate that RAP-SAM achieves the best speed and accuracy trade-off in the proposed benchmark and regular real-time semantic and panoptic segmentation benchmarks. We also show scalability across datasets and application demos.
2 Related Work
Universal Segmentation. Prior research has focused on designing segmentation models tailored for specific tasks. Recent developments in this field [13, 52, 2, 93, 10, 85, 49, 39, 17, 86, 42, 43, 45, 87] have shifted towards employing mask classification architectures, coupled with an end-to-end set prediction objective, to facilitate universal segmentation. This approach has outperformed specialized models [4, 33, 20, 48, 27, 88, 96, 7, 40, 41, 6] across various segmentation tasks, including those involving images, videos, and point clouds [32, 47, 46, 9, 67, 78, 79]. Specifically, Mask2Former [10] uses a masked-attention mechanism and surpasses the performance of prior specific image segmentation models. Similarly, Tube-Link [46] outperforms earlier models [47, 40, 51] tailored for specific video segmentation tasks. Semantic-SAM [35], built on Mask-DINO [36], employs a hybrid query design to facilitate both panoptic and interactive segmentation within a single framework. However, despite the strong performance of these works, their complexity, stemming from the heavier encoder and the cascaded decoder designs, poses significant challenges for integration into actual products. These are primarily hindered by slow processing speeds and substantial parameter overhead.
Vision Foundation Model. Recent advancements in the vision research community have been propelled by the development of Vision Foundation Models (VFMs), including pure-visual pre-training [19], vision-language pre-training [15, 37], and multi-modal models employing visual prompting [75, 1]. SAM [34] stands as a seminal work targeting general interactive segmentation. Subsequent studies [3, 12, 77] have adapted SAM for diverse applications, predominantly employing it as a segmentation tool. However, these Vision Foundation Models typically employ substantial visual encoders, like ViT-base [13], which poses challenges for deployment on real-time or mobile devices due to their size.
Efficient Model Design. This research direction primarily concentrates on the development of efficient CNNs [24, 66, 23, 28, 94, 56, 18], transformers [57, 64, 58], and hybrid architectures [90, 55, 8, 38, 100], aimed at advancing visual representation learning. Most research investigates backbone design within the constraints of real-time processing and parameter efficiency. These studies are orthogonal to our own, which primarily explores the impact of data, task, and decoder design. Additionally, our work employs efficient models as encoders and presents detailed benchmarks and analyses of the encoder’s effects.
Efficient Segmentation. Previous studies [95, 44, 84, 25, 22, 71, 83, 59] on efficient segmentation have predominantly concentrated on closed-set and specific domains. In particular, a significant portion of this research [44, 25, 60] is dedicated to driving scenarios. Recently, various studies [92, 71, 99] have developed efficient segmentation techniques that facilitate model execution on mobile devices. Mobile SAM [89] introduces a streamlined encoder distillation method. Fast SAM [97] employs a single-stage instance segmentation framework that directly decodes class-agnostic masks. Recently, multiple studies have been conducted on efficient panoptic segmentation [69, 25, 63] and rapid video instance segmentation [82, 91]. However, these real-time segmentation methods are limited to specific tasks. We contend that an all-purpose model capable of real-time performance would have wide-ranging applications in editing, tracking, and segmentation functionalities within various products.
3 Problem Setup
Settings. Our proposed real-time all-purpose segmentation contains three visual inputs, including image, video, and visual prompts (boxes and points). It outputs corresponding masks, labels, and instance ID. We adopt panoptic segmentation for image segmentation, which unifies semantic and instance segmentation. For video segmentation, we adopt video instance segmentation, where we aim to segment and track each foreground instance. For interactive segmentation, we follow the SAM-like setting, where we perform class-agnostic segmentation according to the user’s inputs. Compared with previous settings, as shown in Tab. 1, we can achieve more flexible input and output, which is more challenging. Combined with interactive segmentation and video segmentation, we can also achieve interactive video segmentation as video object segmentation.
Datasets. For image segmentation and interactive segmentation, we use the well-known COCO dataset [50] for benchmarking. We adopt partially SAM data [34] for testing. Due to the limited resources and limited model capacity in real-time settings, we do not use SAM data for training. We take COCO data as SAM data for training and testing, where the boxes and random points sampled from masks are used for visual prompting. For video segmentation, we adopt the widely used YouTube-VIS 2019 dataset [81] for training. We do not use the SAM data for our benchmarking. There are several reasons: (1), SAM data is not well annotated as the COCO dataset. (2), We only focus on object-level and scene-level segmentation, while SAM data contain more granularity. (3), SAM data has no semantics, while our goal is to predict mask labels for real applications.
Figure 2: RAP SAM overview. Our method contains three visual inputs: image, video, and visual prompts. Utilizing positional encoding, we generate prompt queries from these visual prompts. The learnable object queries, alongside the prompt queries and the feature map $F$, are directed to the multi-stage decoder. This process generates multi-stage predictions and refined queries. These refined queries engage in cross-attention with $F$, resulting in the final prediction.
4 Proposed Method
Motivation. Employing the same formulation, we first revisit three tasks: panoptic segmentation, video instance segmentation, and interactive segmentation. When combined with a query-based mask transformer, object queries can represent all entities within the scene, encompassing image-level objects, video-level tracked objects, and prompt-based specific objects. Each task’s formulation is as follows:
For panoptic segmentation [33], given an input image $I\in\mathbb{R}^{H\times{W}\times 3}$, the objective is to generate a group of masks $\{y_{i}\}_{i=1}^{N}=\{(m_{i},c_{i})\}_{i=1}^{N}\,$ where $c_{i}$ denotes the ground truth class label of the binary mask $m_{i}$ and $N$ is the number of masks, ${H}\times{W}$ indicates the spatial dimensions.
For video instance segmentation [81], given a video clip input $V\in\mathbb{R}^{T\times H\times{W}\times 3}$, where $T$ denotes the number of frames, the process aims to generate an instance mask tube $\{y_{i}\}_{i=1}^{N}=\{(m_{i},c_{i},d_{i})\}_{i=1}^{N}\,$, where $N$ represents the number of the tube masks $m_{i}\in{\{0,1\}}^{{T}\times{H}\times{W}}$, $c_{i}$ indicates the class label of the tube $m_{i}$, and $d_{i}$ identifies the instance ID for each tube.
@@ 4 Proposed Method
Table 1: Real-Time Segmentation Settings. Our all-purpose segmentation supports more tasks in one framework.
@@
Property Image Seg Video Seg SAM-Like Ours
@@
Table 2: Comparison of Segmentation Methods. Our proposed RAP-SAM supports various segmentation tasks, and it can run in real time.
@@
ICNet [95]
Bi-Seg [84]
YOSO [25]
Mobilie-VIS [91]
SAM [34]
Mask2Former [10]
Video K-Net [47]
OneFormer [29]
RAP-SAM (Ours)
@@
For SAM-like interactive segmentation [34], both the image $I$ and visual prompts $P\in\mathbb{R}^{K\times 4}$ (box prompts are used here) are taken as inputs. These prompts, which include points and boxes, lead to the outputs of corresponding binary image masks $\{y_{i}\}_{i=1}^{K}=\{m_{i}\in\{0,1\}^{H\times W}\}_{i=1}^{K}$, where $K$ denotes the number of visual prompts. Each visual prompt is encoded into one object query, which naturally can be the input of the decoder, as implemented in  [34, 10].
Single Object Query For All Segmentation. As mentioned above, by combining all three settings into one, we can represent all the output segmentation entities using the same query-based mask classification framework. In particular, one object query corresponds to one mask $m_{i}$, one label $c_{i}$, and one ID $d_{i}$. We aim to design a simple, efficient, and parameter-sharing framework to support these tasks.
Compared Baselines. In Sec. 5.1, we benchmark several baselines for our all-purpose segmentation. We use the prompt encoder (proposed in SAM [34]) to encode the visual prompt and jointly learn with object query for both PS and VIS on a shared decoder. Although we adopt the lightweight design to speed up, their performance is still limited or not balanced, as shown in Tab 3.
4.1 RAP-SAM Architecture
@@
Overall Architecture. Our RAP-SAM is a simple encoder and decoder architecture. As shown in Fig. 2, it contains a backbone, a lightweight neck, and a shared multitask decoder. Visual prompts $P$ are also the input of the decoder. Following SAM, we also adopt the prompt encoder to encode visual prompts into a query. We adopt the same decoder for both visual prompts and initial object queries to share more computation and parameters. However, the goals of both are different. The former pays more attention to the local details, while the latter also considers the scene and temporal features. To better balance the results for interactive segmentation and image/video segmentation, we design a prompt adapter and an object adapter in the end of the decoder.
@@
Lite Feature Extractor. Due to computation cost limitation, we avoid the design of previous methods, such as a large backbone and heavier transformer encoder. We explore lightweight backbones, including ResNet18 [21], STDC-v1 [14], and SeaFormer [71]. We adopt feature pyramid networks with deformable convolutions to fuse multi-scale features and obtain more aligned feature representation, motivated by previous works [44] on real-time semantic segmentation. For video input, we extract the spatial-temporal features. For simplicity, we use $F_{img}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times d}$ for image inputs and $F_{vid}\in\mathbb{R}^{T\times\frac{H}{4}\times\frac{W}{4}\times d}$ for video inputs.
@@
Unified Dynamic Convolution Decoder. As in previous works, the goal of the decoder is to refine the object query. However, many of these approaches are impractical for real-time settings due to their reliance on heavily cascaded layers and pixel-wise cross-attention mechanisms. In contrast, our method employs a pooling-based dynamic convolution framework [93, 47] to enhance the decoder’s efficiency.
Given object query $Q_{i}\in\mathbb{R}^{N\times d}$ and prompt query $P_{i}\in\mathbb{R}^{K\times d}$, the initial masks $M_{i}$ are obtained by dot product with $F_{img}$ or $F_{vid}$. $i$ denotes the index of the decoder stage. According to the inputs, the masks can be image masks for panoptic segmentation $M_{i}^{pan}\in\mathbb{R}^{N\times\frac{H}{4}\times\frac{W}{4}}$ and interactive segmentation $M_{i}^{iter}\in\mathbb{R}^{K\times\frac{H}{4}\times\frac{W}{4}}$ or tube masks $M_{i}^{tube}\in\mathbb{R}^{N\times T\times\frac{H}{4}\times\frac{W}{4}}$. Then, we can obtain the corresponding query features $X_{i}^{pan}\in\mathbb{R}^{N\times d}$, $X_{i}^{iter}\in\mathbb{R}^{K\times d}$, $X_{i}^{tube}\in\mathbb{R}^{N\times d}$ via mask pooling:
@@
In particular, we adopt the design [93, 68, 42] to refine input queries $Q_{i}^{pan}$ with features $X_{i}^{pan}$ which are grouped from their masks,
@@
where the dynamic convolution uses the query features $X^{pan}_{i}$ to generate parameters to weight input queries $Q_{i-1}^{pan}$. Specifically, $\mathrm{DynamicConv}$ uses input query features $X^{pan}_{i}$ to generate gating parameters via a multilayer perceptron (MLP) and multiply back to the original query input $Q^{pan}_{i-1}$. We adopt the same design [93, 68] by learning gating functions to update the refined queries. The $\mathrm{DynamicConv}$ operation is based on:
@@
where $\mathrm{Gate}$ is implemented with a fully connected (FC) layer followed by Layer Norm (LN) and a sigmoid layer. We adopt two different gate functions, including $\mathrm{Gate}_{x}$ and $\mathrm{Gate}_{q}$. They are implemented by MLP. The former is to weigh the query features, while the latter is to weigh corresponding queries. Next, we adopt one self-attention layer with feed-forward layers [70, 72] to learn the correspondence among each query and update them accordingly. This operation leads to the full correlation among queries, shown as follows:
@@
where $\mathrm{MHSA}$ means Multi Head Self Attention, $\mathrm{FFN}$ is the Feed Forward Network commonly used in the current version of Transformers [2, 13]. Finally, the refined masks are obtained via dot product between the refined queries $Q^{pan}_{i}$, and the input features $F_{p}$, $F_{s}$. For mask classification, we adopt several feed-forward layers on $Q^{pan}_{i}$ and directly output the class scores. Then, we perform the inner product between the learned queries $Q^{pan}_{i}$, $Q^{iter}_{i}$, and their features $F_{vid}$ and $F_{img}$. For the video segmentation task, the only difference is that we predict the tube mask rather than the image mask. The entire process (Equ. 1, Equ. 2, Equ. 3 and Equ. 5) is repeated three times in total.

### v2

# RMP-SAM: Towards Real-Time Multi-Purpose Segment Anything

@@ Abstract
Recent segmentation methods, which adopt large-scale data training and transformer architecture, aim to create one foundation model that can perform multiple tasks. However, most of these methods rely on heavy encoder and decoder frameworks, hindering their performance in real-time scenarios. To explore real-time segmentation, recent advancements primarily focus on semantic segmentation within specific environments, such as autonomous driving. However, they often overlook the generalization ability of these models across diverse scenarios. Therefore, to fill this gap, this work explores a novel real-time segmentation setting called real-time multi-purpose segmentation. It contains three fundamental sub-tasks: interactive segmentation, panoptic segmentation, and video instance segmentation. Unlike previous methods, which use a specific design for each task, we aim to use only a single end-to-end model to accomplish all these tasks in real-time. To meet real-time requirements and balance multi-task learning, we present a novel dynamic convolution-based method, Real-Time Multi-Purpose SAM (RMP-SAM). It contains an efficient encoder and an efficient decoupled adapter to perform prompt-driven decoding. Moreover, we further explore different training strategies and one new adapter design to boost co-training performance further. We benchmark several strong baselines by extending existing works to support our multi-purpose segmentation. Extensive experiments demonstrate that RMP-SAM is effective and generalizes well on proposed benchmarks and other specific semantic tasks. Our implementation of RMP-SAM achieves the optimal balance between accuracy and speed for these tasks. The code is released at https://github.com/xushilin1/RAP-SAM
@@ 1 Introduction
Recent advancements in computer vision, driven by transformer architectures (Dosovitskiy et al., 2021; Carion et al., 2020; Wang et al., 2023b), focus on creating a single large-scale model versatile enough to handle different tasks like detection and segmentation with a general-purpose design. Several studies utilizing the same architectures (Wang et al., 2023a; Yan et al., 2023; Jain et al., 2023) have shown superior performance across various tasks. Enjoying the benefit of the large-scale model capacity and training data, the general-purpose models demonstrate strong performance on versatility for specific applications (Kirillov et al., 2023; Jia et al., 2021). For image and video segmentation at the semantic level, numerous studies (Cheng et al., 2022; Zhang et al., 2021; Li et al., 2022d; 2023e; Jain et al., 2023) employing a universal design have surpassed previous models tailored for specific image and video segmentation tasks. The Segment Anything Model (SAM) (Kirillov et al., 2023) is a leading approach in interactive segmentation, offering a unified model for various interactions such as point, bounding box, mask, and text input.
Figure 1: We present real-time multi-purpose segmentation to segment and recognize objects for image, video, and visual prompt inputs. In addition to benchmarking, we propose a simple yet effective baseline, named RMP-SAM, which achieves the best performance and speed trade-off among three different tasks. The larger dot indicates more parameters.
However, most of these studies face challenges in achieving real-time performance on devices with limited computing resources due to their heavy encoders or cascaded decoders. Despite their high-performance capabilities and versatility, this limitation hinders the practical use of these advanced architectures in vision applications, such as editing tools on edge devices. In contrast, several studies (Zhao et al., 2018; Hu et al., 2023) focus on real-time segmentation using specific designs (e.g., efficient decoder fusion) but are often tailored to a specific application trained on a particular dataset (e.g., COCO) or restricted to similar tasks like video semantic segmentation. To our knowledge, no study investigates real-time, multi-purpose segmentation, where a single model can perform universal segmentation tasks, including image, video, and SAM-like interactive segmentation. For example, given a video input, we want to create a tool to perform various segmentation tasks on the device, including interactive segmentation with visual cues, labeling the masks, and tracking the masks.
@@
Inspired by the abovementioned challenges, we focus on real-time multi-purpose segmentation and efficiently investigate the novel issue of training versatile models. We pose a critical inquiry: Given the limited computational resources and model capacity, how can we develop an efficient, multi-purpose segmentation model? This entails creating a single model capable of segmenting, tracking, and classifying each pixel in real-time, similar to performing interactive segmentation similar to SAM. We present an illustration figure in Fig. 1(a).
@@
First, we explore this problem by benchmarking existing methods. Specifically, we use the COCO (Lin et al., 2014) and YouTube-VIS (Yang et al., 2019) for joint co-training, employing identical hyperparameters. Beyond the object queries used for image and video segmentation in prior research, our approach incorporates visual prompt queries for interactive segmentation. However, none of these works can achieve the best accuracy and speed trade-off on three tasks using one model. This is because: 1) No previous works explore joint co-training for segmentation for different purposes. Thus, extensive experiments are needed to gain a better understanding of meta-architecture. 2) Interactive and semantic-level segmentation needs better strategies to balance each other. 3) Under the real-time setting, several previous architectures (Yan et al., 2023; Li et al., 2022c) for multi-task learning may not be practical.
@@
To solve these issues, we present a real-time multi-purpose segment anything model (RMP-SAM). It contains an efficient encoder and shared decoder with multiple queries as inputs. Unlike previous works, we avoid cascaded transformers encoder layers and pixel-wise cross-attention decoder (Cheng et al., 2022). Instead, we propose a pooling-based dynamic convolution to replace per-pixel cross-attention to achieve better accuracy and speed trade-off. We represent all three tasks using queries with shared dynamic convolution operation. According to different outputs, we can divide the segmentation tasks into two types: one for class-agnostic masks (SAM-like) and the other for class-specific masks (contains one mask and one class label). Two types have different goals and will conflict during the multi-task learning process. Thus, we further explore four different architectures in the case of various decoder designs. In particular, we find a single decoder with multiple queries for different tasks can achieve good enough results and best speed and accuracy trade-off. Moreover, we present an effective dual adapter design that better adapts shared knowledge of the same decoder. The dual adapter contains an asymmetric design to better transfer shared features. To the best of our knowledge, RMP-SAM is the first real-time multi-purpose segmentation model that can achieve the best speed and accuracy trade-off on four different tasks, as shown in Fig. 1(b). Moreover, due to the semantic level and interactive level segmentation ability, our model can create more new applications, such as interactive video segmentation. In addition, we also verify our model in the SAM-like setting by co-training with SAM-data, it also achieves comparable results. Extensive experiments show the effectiveness of our design. The main contributions of this work are:
@@
  • We introduce real-time multi-purpose segmentation, a novel multi-task segmentation aiming to segment objects for real-time image, video, and interactive inputs.
@@
  • We present a simple yet fast baseline named RMP-SAM. It contains a lightweight feature extractor, a unified decoder, and two asymmetric adapters.
@@
  • Extensive experiments demonstrate that RMP-SAM achieves the best speed and accuracy trade-off in the proposed benchmark and regular real-time semantic and panoptic segmentation benchmarks. We also show scalability across datasets and application demos.
2 Real-Time Multi-Purpose Segmentation
@@
Table 1: Real-Time Segmentation Settings. Our multi-purpose segmentation supports more tasks in one framework.
@@
Property Image Seg Video Seg SAM-Like SAM-2-Like Ours: Real-Time Multi-Purpose Seg
@@
Table 2: Comparison of Segmentation Methods. Our proposed RMP-SAM supports various segmentation tasks and performs them in real-time.
@@
ICNet (Zhao et al., 2018)
Bi-Seg (Yu et al., 2018)
YOSO (Hu et al., 2023)
Mobilie-VIS (Zhang et al., 2024)
SAM (Kirillov et al., 2023)
Mask2Former (Cheng et al., 2022)
Video K-Net (Li et al., 2022d)
OneFormer (Jain et al., 2023)
RMP-SAM (Ours)
Problem Settings. Our proposed real-time multi-purpose segmentation contains three visual inputs: image, video, and visual prompts (boxes and points). It predicts the corresponding masks, labels, and instance IDs for the content within the given input. Basically, we adopt the definition of panoptic segmentation Kirillov et al. (2019) for image inputs since it covers all concepts in the scene. For video inputs, we segment and track each foreground instance (video instance segmentation). Given visual prompts, we follow the SAM-like Zhang et al. (2023a); Zhou et al. (2023); Shu et al. (2023); Xiong et al. (2023) setting for interactive segmentation, performing class-agnostic segmentation according to the user’s inputs. Although recent advancements, such as SAM-2 (Ravi et al., 2024), have attempted to unify the interactive segmentation task for both image and video inputs, our framework emphasizes the importance of object semantics. It incorporates multi-task learning in a single shot. Combined with interactive segmentation and video segmentation, we can also achieve interactive video segmentation as video object segmentation. As shown in Tab. 1 and Tab. 2, we compare previous segmentation methods with and indicate that our method unifies different segmentation tasks within a single framework. This approach enables more flexible usage across various segmentation tasks, particularly regarding input modalities and predictions.
Datasets. For benchmarking, we use the well-known COCO dataset (Lin et al., 2014) for panoptic and interactive segmentation. During the training of interactive segmentation, the boxes derived from the ground truth, along with points randomly sampled from the masks, serve as visual prompts. For testing, we obtain the center point of the mask as a visual prompt. For video segmentation, we adopt the widely used YouTube-VIS 2019 dataset (Yang et al., 2019) for training. We do not use the SAM data for our benchmarking. There are several reasons: (1) SAM data is not as well annotated as the COCO dataset. (2) We only focus on object-level and scene-level segmentation, while SAM data contain more granularity. (3) SAM data has no semantics, while our goal is to predict semantic labels for real applications. To fairly compare with SAM-like models, we adopt the SAM-data training for specific setting, as shown in Sec. 4.1.
3 Proposed Method
@@
First, we revise panoptic segmentation, video instance segmentation, and interactive segmentation under the unified formulation. When combined with a query-based segmentation model, object queries can represent all entities within the scene, encompassing image-level objects, video-level tracked objects, and prompt-based specific objects. For panoptic segmentation (Kirillov et al., 2019), given an input image $I\in\mathbb{R}^{H\times{W}\times 3}$, the objective is to generate a group of masks $\{y_{i}\}_{i=1}^{N}=\{(m_{i},c_{i})\}_{i=1}^{N}\,$ where $c_{i}$ denotes the class label of the binary mask $m_{i}$ and $N$ is the number of queries, ${H}\times{W}$ indicates the spatial dimensions. For video instance segmentation (Yang et al., 2019), given a video clip input $V\in\mathbb{R}^{T\times H\times{W}\times 3}$, where $T$ denotes the number of frames, the process aims to generate an instance mask $\{y_{i}\}_{i=1}^{N}=\{(m_{i},c_{i},d_{i})\}_{i=1}^{N}\,$, where $N$ represents the number of the tube masks $m_{i}\in{\{0,1\}}^{{T}\times{H}\times{W}}$, $c_{i}$ indicates the class label of the tube $m_{i}$, and $d_{i}$ identifies the instance ID for each instance. For SAM-like segmentation (Kirillov et al., 2023), both the image $I$ and visual prompts $P\in\mathbb{R}^{K\times 2}$ (point prompts are used here) are taken as inputs. These prompts, which include points and boxes, lead to the outputs of corresponding binary image masks $\{y_{i}\}_{i=1}^{K}=\{m_{i}\in\{0,1\}^{H\times W}\}_{i=1}^{K}$, where $K$ denotes the number of visual prompts. Each visual prompt is encoded into one object query, which naturally can be the input of the decoder.
3.1 RMP-SAM Architecture
@@
Overall Architecture. As shown in Fig. 2, our RMP-SAM is built upon a simple encoder and decoder architecture that can input images, videos, and visual prompts. Following SAM, we utilize the prompt encoder to convert visual prompts into queries. These visual prompts and object queries share the same decoder for the efficiency of computation and parameter size. However, the goals of the two types of queries are different: the former pays more attention to the local details of the visual prompts, while the latter considers the scene and temporal features. To better balance the results for interactive segmentation and image/video segmentation, we design a prompt adapter and an object adapter at the end of the decoder.
Figure 2: RMP-SAM overview. Our method contains three visual inputs: image, video, and visual prompts. Utilizing positional encoding, we generate prompt queries from these visual prompts. The learnable object queries, prompt queries, and the feature map $F$ are directed to the multi-stage decoder. This process generates multi-stage predictions and refined queries. These refined queries engage in cross-attention with $F$, resulting in the final prediction.
@@
Lite Feature Extractor. Due to computation cost limitations, we avoid using previous methods (Cheng et al., 2022; Li et al., 2024b) that involve large backbones and heavier encoders. As required by real-time constraint, we investigate lightweight backbones such as ResNet18 (He et al., 2016) and TopFormer (Zhang et al., 2022). We incorporate a feature pyramid network with deformable convolution to enhance representation alignment and fuse multi-scale features in real-time. When dealing with video input, we extract the spatial-temporal features. For simplicity, we use $F_{img}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times d}$ for image inputs and $F_{vid}\in\mathbb{R}^{T\times\frac{H}{4}\times\frac{W}{4}\times d}$ for video inputs.
@@
Unified Dynamic Convolution Decoder. The goal of the decoder is to refine the object query. However, many of these methods (Cheng et al., 2021b; 2022), are not feasible for real-time settings due to their reliance on heavily cascaded layers and pixel-wise cross-attention mechanisms. In contrast, our method uses a pooling-based dynamic convolution framework (Zhang et al., 2021; Li et al., 2022d) to enhance the efficiency of the decoder. Given object query $Q\in\mathbb{R}^{N\times d}$ and prompt query $P\in\mathbb{R}^{K\times d}$, the masks $M_{i}$ are obtained by dot product with $F_{img}$ or $F_{vid}$. Here, $i$ represents the decoder stage index, $N$ is the total number of object queries, and $K$ is the number of visual prompts. According to the inputs, the masks can be image masks for panoptic segmentation $M_{i}^{pan}\in\mathbb{R}^{N\times\frac{H}{4}\times\frac{W}{4}}$ and interactive segmentation $M_{i}^{iter}\in\mathbb{R}^{K\times\frac{H}{4}\times\frac{W}{4}}$ or tube masks $M_{i}^{tube}\in\mathbb{R}^{N\times T\times\frac{H}{4}\times\frac{W}{4}}$. Then, we can obtain the corresponding query features $X_{i}^{pan}\in\mathbb{R}^{N\times d}$, $X_{i}^{iter}\in\mathbb{R}^{K\times d}$, $X_{i}^{tube}\in\mathbb{R}^{N\times d}$ via mask pooling:
@@
In particular, we adopt the design of previous methods (Zhang et al., 2021; Sun et al., 2021; Li et al., 2022c) to refine input queries $Q_{i}^{pan}$ with features $X_{i}^{pan}$ which are grouped from their masks,
@@
where the dynamic convolution uses the query features $X^{pan}_{i}$ to generate parameters to weight input queries $Q_{i-1}^{pan}$. Specifically, $\mathrm{DynamicConv}$ uses input query features $X^{pan}_{i}$ to generate gating parameters via a multilayer perceptron (MLP) and multiply back to the original query input $Q^{pan}_{i-1}$. We adopt the same design (Zhang et al., 2021; Sun et al., 2021) by learning gating functions to update the refined queries. The $\mathrm{DynamicConv}$ operation is based on:
@@
where $\mathrm{Gate}$ is implemented with a fully connected layer followed by Layer Norm and a sigmoid layer. We adopt two different gate functions, including $\mathrm{Gate}_{x}$ and $\mathrm{Gate}_{q}$. The former is to weigh the query features, while the latter is to weigh corresponding queries. Next, we adopt one self-attention layer with feed-forward layers (Vaswani et al., 2017; Wang et al., 2021) to learn the correspondence among each query and update them accordingly. This operation leads to the full correlation among queries, shown as follows:
@@
where $\mathrm{MHSA}$ means Multi Head Self Attention, $\mathrm{FFN}$ is the Feed Forward Network commonly used in the current version of Transformers (Carion et al., 2020; Dosovitskiy et al., 2021). We adopt several feed-forward layers on $Q^{pan}_{i}$ for mask classification and directly output the class scores. Then, we perform the inner product between the learned queries $Q^{pan}_{i}$, and image features $F_{vid}$ and $F_{img}$. The only difference for the video segmentation task is that we predict the tube mask rather than only a single image mask. The entire process (Equ. 1, Equ. 2, Equ. 3 and Equ. 5) is repeated three times in total.